OpenStudy (anonymous):

y=1/x-2,x=4 the slope of the given curve at x=4 is?

OpenStudy (tkhunny):

Can you find the 1st Derivative?

OpenStudy (anonymous):

i think its -1/(x-2)^2

OpenStudy (anonymous):

not sure though

OpenStudy (anonymous):

It's all fine... now just substitute x=4 in here...

OpenStudy (raden):

it means : m = -1/(x-2)^2 = -1/(4-2)^2 = ........

OpenStudy (tkhunny):

BTW, you wrote the original problem incorrectly. What you wrote is \(\dfrac{1}{x} - 2\). This seems not to be what was intended.